# SyncLipMAE: Contrastive Masked Pretraining for Audio–Visual Talking-Face Representations

### v1

@@ 5 Experiments
where larger values indicate a clearer minimum at the correct offset (stronger A/V synchrony) Chung & Zisserman (2016); Prajwal et al. (2020). FID measures per-frame perceptual quality via the Fréchet distance between Inception features of real vs. generated frames (lower is better) Heusel et al. (2017); FVD extends this to spatiotemporal video features to capture visual quality and temporal coherence jointly (lower is better) Unterthiner et al. (2018).
@@ Results.
Qualitatively (Fig. 2), SyncLipMAE preserves identity and fine lip details (e.g., teeth, beard) and is more robust to occlusions than LatentSync/MuseTalk. Quantitatively (Tab. 4), consistent with these observations, the audio-driven variant (A-SyncLipMAE) attains the best overall scores on synchrony ($\mathrm{Sync}_{\mathrm{conf}}$) and perceptual quality (FID/FVD) on both VFHQ and Hallo3. The video-driven variant (V-SyncLipMAE) closely matches the audio-driven performance and surpasses prior methods, indicating that the shared token space indeed aligns modalities and enables no-difference audio- or video-driven control within a single model.
@@
We perform ablations on the key components of SyncLipMAE, examining: (i) whether to enable Two-Bypass Face-Aware Masking; (ii) the strategy for adapting audio features; and (iii) the inclusion of cross-attention in the decoder. Detailed results are provided in the Appendix.
@@ 6 Conclusion
Appendix A Ablation Studies
@@
Variant K = 1 K = 5 K = 15 R-precision (32) Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ (A) Two-Bypass Face-Aware Masking A1  Uniform-only (single view) 9.34 7.18 12.68 6.22 18.57 5.36 2.83 5.14 7.92 A2  Face-aware-only (single view) 45.63 3.45 61.07 2.37 75.88 1.27 33.45 52.81 66.19 A3  Two-bypass w/o photometric 49.62 3.02 65.91 2.03 79.87 1.05 37.54 57.89 70.86 (B) Audio Feature Adaptation B1  Last-layer only 42.68 3.62 58.34 2.51 73.45 1.36 30.92 49.87 63.11 (C) Decoder Cross-Attention C1  No cross-attention 48.27 3.10 64.31 2.09 78.44 1.09 36.21 56.48 69.32 Ours 52.53 2.66 68.41 1.73 82.27 0.93 40.18 61.48 73.49
@@
Appendix B More Results
@@
Experimental analysis. Across all four datasets, SyncLipMAE shows consistent gains in both temporal alignment (higher $\mathrm{Acc}_{\pm K}$, lower Offset) and cross-modal retrieval (higher R-precision@k). Improvements are most pronounced on the controlled emotion corpora (MEAD/RAVDESS), where reduced acoustic and visual variability amplifies the benefit of our token-level alignment; on the in-the-wild CelebV-HQ/CelebV-Text, gains remain steady despite non-speech spans and repeated content. Notably, all results are achieved with single-frame visual conditioning (n=1n{=}1), indicating strong per-frame audio–visual correspondence without reliance on long temporal windows.
Appendix C Limitation

### v2

@@ 5 Experiments
where larger values indicate a clearer minimum at the correct offset (stronger A/V synchrony). FID measures per-frame perceptual quality via the Fréchet distance between Inception features of real and generated frames (lower is better) (fid), and FVD extends this to spatiotemporal video features to jointly capture visual quality and temporal coherence (lower is better) (fvd).
@@ Results.
Qualitatively (Fig. 4), SyncLipMAE preserves identity and fine lip details (e.g., teeth, beard) and is more robust to occlusions than LatentSync/MuseTalk. Quantitatively (Tab. 5), consistent with these observations, the audio-driven variant (A-SyncLipMAE) attains the best overall scores on synchrony ($\mathrm{Sync}_{\mathrm{conf}}$) and perceptual quality (FID/FVD) on VFHQ and Hallo3. The video-driven variant (V-SyncLipMAE) closely matches the audio-driven performance and surpasses prior methods, indicating that the shared token space indeed aligns modalities and enables no-difference audio- or video-driven control within a single model.
@@
We perform ablations on the key components of SyncLipMAE, examining: (i) whether to enable Two-Bypass Face-Aware Masking; (ii) the strategy for adapting audio features; (iii) the inclusion of cross-attention in the decoder; and (iv) the composition of the prompt tokens. Detailed results are provided in the Appendix.
@@ 6 Conclusion
Table 7: Statistics of the talking-face pretraining corpus and the downstream tasks that each dataset is used for. Durations are reported in hours.
Dataset #Clips Duration (h) Pretrain (AV sync) Emotion / action cls. VSR Video dubbing HDTF 5,433 14.51 $\checkmark$ $\times$ $\checkmark$ $\checkmark$ MEAD 203,013 243.78 $\checkmark$ $\times$ $\times$ $\checkmark$ RAVDNESS 1,440 1.51 $\checkmark$ $\times$ $\times$ $\checkmark$ CelebV-HQ 35,662 68.07 $\checkmark$ $\checkmark$ $\times$ $\checkmark$ CelebV-Text 67,005 261.23 $\checkmark$ $\times$ $\times$ $\checkmark$ Hallo3 100,229 132.03 $\checkmark$ $\times$ $\checkmark$ $\checkmark$ VFHQ 15,174 36.94 $\checkmark$ $\times$ $\checkmark$ $\checkmark$ Total 427,956 758.07
Appendix A Dataset Statistics
Our pretraining corpus aggregates 427,956427{,}956 talking-face clips spanning approximately $758$ hours of video from seven public datasets: HDTF, MEAD, RAVDESS, CelebV-HQ, CelebV-Text, Hallo3, and VFHQ. As summarized in Tab 7, HDTF, MEAD, RAVDESS, CelebV-HQ, CelebV-Text, Hallo3, and VFHQ contribute $14.51$, $243.78$, $1.51$, $68.07$, $261.23$, $132.03$, and $36.94$ hours respectively. All seven datasets are used for our self-supervised pretraining (AV sync), while different subsets are activated for downstream evaluation: CelebV-HQ is the only source for emotion/action classification, HDTF/Hallo3/VFHQ support VSR benchmarks, and all datasets are used for video dubbing.
The datasets differ markedly in acquisition conditions and semantic coverage. HDTF mainly contains high-definition frontal talking-head videos of news-like broadcasts, offering clean lip motion and relatively simple backgrounds that are well-suited for lip synchronization and VSR. MEAD and RAVDESS are studio-recorded emotional corpora where multiple actors read fixed sentences under controlled lighting and camera setups with discrete emotion labels; MEAD further provides multi-view recordings from several camera angles, while RAVDESS focuses on carefully validated emotional speech and song. CelebV-HQ and CelebV-Text are large-scale, in-the-wild celebrity video datasets with rich annotations (attributes or paired texts) and diverse appearance, head poses, and actions; however, the subjects are not always strictly “talking”—clips may include singing, non-verbal expressions, or other activities—so we primarily rely on them for pretraining and video dubbing rather than VSR supervision. Hallo3 provides high-quality, highly dynamic portrait videos with varied camera viewpoints and rich real-world scenes (e.g., interviews, TV dramas, conversational clips), where audio can contain background sounds in addition to speech, making it particularly suitable for stress-testing robustness in VSR and dubbing. VFHQ is a high-fidelity video face dataset originally curated for video face super-resolution and consists mainly of diverse interview-style talking heads, including singing and non-English speech, which we repurpose as another strong source of high-quality portrait footage for pretraining, VSR, and dubbing.
Overall, combining these sources yields a corpus that balances scale and diversity across identities, poses, emotions, recording environments, and linguistic content. This mix is crucial for learning synchronization-aware facial dynamics that transfer reliably from controlled studio-style datasets (MEAD/RAVDESS) to in-the-wild, multi-scene talking-head footage (CelebV-HQ, CelebV-Text, Hallo3, VFHQ), and for supporting a broad suite of downstream tasks spanning AV synchronization, emotion/action classification, VSR, and video dubbing.
Appendix B Ablation Studies
@@
Variant K = 1 K = 5 K = 15 R-precision (32) Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ (A) Two-Bypass Face-Aware Masking A1  Uniform-only (single view) 9.34 7.18 12.68 6.22 18.57 5.36 2.83 5.14 7.92 A2  Face-aware-only (single view) 45.63 3.45 61.07 2.37 75.88 1.27 33.45 52.81 66.19 A3  Two-bypass w/o photometric 49.62 3.02 65.91 2.03 79.87 1.05 37.54 57.89 70.86 (B) Audio Feature Adaptation B1  Last-layer only 42.68 3.62 58.34 2.51 73.45 1.36 30.92 49.87 63.11 (C) Decoder Cross-Attention C1  No cross-attention 48.27 3.10 64.31 2.09 78.44 1.09 36.21 56.48 69.32 (D) Prompt Tokens D1  $z^{\mathrm{voc}}$ 47.12 3.18 63.27 2.16 77.95 1.12 35.64 55.73 68.41 D2  $z^{\mathrm{id}},z^{\mathrm{voc}}$ 50.03 2.90 66.47 1.96 80.56 1.00 38.22 59.03 71.92 D3  $z^{\mathrm{amb}},z^{\mathrm{voc}}$ 51.41 2.74 67.39 1.82 81.33 0.97 39.26 60.37 72.81 Ours; $z^{\mathrm{id}},z^{\mathrm{voc}},z^{\mathrm{amb}}$ 52.53 2.66 68.41 1.73 82.27 0.93 40.18 61.48 73.49
Table 9: Lip synchronisation on HDTF.
Method K = 1 K = 5 K = 15 R-precision (32) Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ SyncNet 37.91 3.91 46.92 2.70 55.10 1.66 20.81 32.92 41.83 VocalLiST -LRS2 40.61 3.57 46.99 2.77 54.94 1.80 24.08 36.04 43.75 StableSyncNet 54.24 1.64 55.51 1.56 57.04 1.46 46.77 62.26 71.11 SyncLipMAE (ours) 44.98 3.04 57.70 2.09 69.67 1.29 43.70 63.52 74.41
@@
The Composition of the Prompt Tokens
As introduced in section 3.1, SyncLipMAE encodes each talking-face frame into a three-token representation $(\mathbf{z}^{\mathrm{id}},\mathbf{z}^{\mathrm{amb}},\mathbf{z}^{\mathrm{voc}})$. In this experiment, we examine the necessity and effectiveness of this design by ablating the prompt-token composition: we remove $\mathbf{z}^{\mathrm{id}}$, $\mathbf{z}^{\mathrm{amb}}$, or both, and then evaluate how these variants affect the quality of the extracted $\mathbf{z}^{\mathrm{voc}}$ on the AV-Sync task. As shown in Experiment D, dropping any of the tokens consistently degrades the ability of $\mathbf{z}^{\mathrm{voc}}$ to match the audio tokens, which we attribute to the orthogonality loss in our objective that explicitly encourages the three factors to focus on complementary aspects of the talking-face frame.
Appendix C More Results
@@
Table 14: Lip synchronisation on out-of-distribution data. Higher is better for Acc and R-Precision; lower is better for Offset.
Method K = 1 K = 5 K = 15 R-precision (32) Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Acc $\uparrow$ (%) Offset $\downarrow$ Top1 $\uparrow$ Top2 $\uparrow$ Top3 $\uparrow$ SyncNet -5 25.98 5.19 33.38 4.15 41.34 3.08 13.31 21.21 27.99 VocaLiST -5 25.77 5.10 29.82 4.51 35.46 3.73 13.28 21.44 27.51 StableSyncNet -16 36.20 3.19 37.53 3.10 40.49 2.89 27.37 38.57 46.50 SyncLipMAE  -1 (ours) 40.30 3.78 50.71 3.02 61.67 2.25 33.66 51.12 61.33
[…]
Zero-shot Comparison
Given that prior state-of-the-art methods and our approach differ in both training procedure and training data sources, it is important to compare them under a truly zero-shot setting on datasets that none of the methods have seen during training. For AV synchronization, we therefore collected 100 public speaking videos from the internet and evaluated all methods in a zero-shot manner on this set. For VSR, neither our model nor the baselines are trained on CelebV-HQ or CelebV-Text, so we additionally constructed a zero-shot test set by selecting 50 clips with normal talking behaviour from these datasets. As shown in  Tab 14 and  Tab 15, SyncLipMAE also outperforms previous state-of-the-art methods on datasets it has never been trained on.
Experimental analysis. We observe that on the AV-Sync task, SyncLipMAE exhibits surprisingly strong performance. Despite operating on single frames only, SyncLipMAE shows consistent gains in both temporal alignment (higher $\mathrm{Acc}_{\pm K}$, lower Offset) and cross-modal retrieval (higher R-precision@K). Thanks to being trained on talking-face datasets that cover a wide range of scenarios, this advantage is particularly pronounced in generic in-the-wild settings and in the presence of environmental noise (Hallo3, VFHQ, zero-shot evaluation). On some datasets (HDTF, MEAD), however, the purely single-frame design yields noisier behavior when we aggregate frame-wise predictions over a short temporal window, leading to slightly weaker temporal alignment than the multi-frame counterpart at K=1K{=}1 and K=5K{=}5. As $K$ increases, the advantages of SyncLipMAE become increasingly pronounced. For VSR, the well-aligned audio–visual representations learned by SyncLipMAE make it easier for the model to infer the underlying speech content from the observed visual cues.
Appendix D Limitation